OpenStudy (anonymous):

.Karen plotted the route of a city tour bus on a coordinate grid. It shows the route as a quadrilateral with vertices at (2, 5), (4, 9), (8, 7), and (6, 3). Which statement about the bus route is true? Answer It is a square of side length 2 times square root of 5 units. It is a square of side length 2 units. It is a rectangle of width 2 units and length 4 units. It is a rectangle of width 2 times square root of 5 units and length 4 times square root of 5 units.

OpenStudy (anonymous):

Same idea. Caculate four distances of four sides using distance formula. Then, the answer will be clear.

OpenStudy (anonymous):

Or, plot on a coordinate graph and see what shape the points form.

OpenStudy (anonymous):

its a square. but i dont understand the first two answer choices

OpenStudy (anonymous):

Side length = 2 or 2*SQRT(5)?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes, i dont really understand it

OpenStudy (anonymous):

Distance between: (2, 5), (4, 9) IS for example: SQRT(16+4) = SQRT(20) = SQRT(4) * SQRT(5) = 2 * SQRT(5)